OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (6, −8) and has a slope of −2?

OpenStudy (anonymous):

y – y1 = m(x – x1) If we plug in x and y, y + 8 = m(x - 6) If we plug in m, y + 8 = -2(x - 6) This is point slope form.

OpenStudy (anonymous):

ty :)